Coves Cakes is a local bakery. Price and cost information follows: $ 14.71 Price per cake Variable cost per cake IngredientsCoves Cakes is a local bakery. Price and cost information follows: $ 14.71 Price per cake Variable cost per cake IngredientsCoves Cakes is a local bakery. Price and cost information follows: Price per cake Variable cost per cake Ingredients Direct
Answer #1

Answer

  • All working forms part of the answer
  • Requitrement 1

A

Price per cake

$                    14.71

B

Total variable cost per cake

$                       3.52

C = A - B

Contribution margin per cake

$                    11.19

D

Fixed Cost

$              5,147.40

E = C/D

Break Even Units

460

Answer

F = E x A

Break Even Sales Dollars

$              6,766.60

Answer

  • Requirement 2

A

Sold Units

540

B

Break Even Units

460

C = A - B

Margin of Safety Units

80

D = C x 14.71]

Margin of Safety

$              1,176.80

Answer

  • Requirement 3

A

Target Profits

$              1,900.00

B

Fixed Cost

$              5,147.40

C = A+B

Total Contribution margin required

$              7,047.40

D

Contribution margin per cake

$                    11.19

E = C/D

Target Sales Units

630

Answer